OpenStudy (anonymous):

show me the solution of these problem if the numerator of a certain fraction is increase by 6 and its denominator is decrease by 5, the resulting fraction is equal to 3/4, if the reciprocal of the original fractionis decrease by 1, the resulting fraction is 16/9, find the origianl fraction?

OpenStudy (anonymous):

x+6/y-5=3/4 (y-5/x+6)=16/9 I think this is how you set it up....

OpenStudy (anonymous):

yea i think misterchyme did good. now you just have to solve the system of equations.

OpenStudy (anonymous):

just show the first step

OpenStudy (anonymous):

oh wait the second equation is (y-5/x+6)-1=16/9

OpenStudy (anonymous):

show me please

OpenStudy (anonymous):

well you would pick which equation looks the easiest to separate variables. I think the second one will actually be best for this.. First you would add one to each side and get (y-5)/(x+6)=25/9 then multiply both sides by x+6 y-5=(x+6)25/9 solve for y in terms of x a then plug it into the other equation.. This one is going to get kinda messy, but should be fun! Can you take it from there?

OpenStudy (anonymous):

im hanging, can u just show me how

OpenStudy (anonymous):

am i going to find the lcd

OpenStudy (anonymous):

hmm im going to have to think, the way i was going about it didnt really give a solution the x's just cancelled out

OpenStudy (anonymous):

WOW I just posted the answer and the work... and it didnt show up.. /fail

OpenStudy (anonymous):

x+6/y-5=3/4 (y-5/x+6)=25/9 OKAY lets try this again... I am going to get rid of the fractions so it looks less intimidating and is easier to work with.. watch carefully 1st equation- Take the reciprocal of 3/4 and multiply the other side by it 4(x+6)/3(y-5)=1 4x+24=3y-15 4x-3y=-39 2nd equation Do the same thing but with 25/9( which is 1+16/9) 9(y-5)/25(x+6)=1 9y-45=25x+150 9y-25x=195 Now we have two equations that look much easier to work with 4x-3y=-39 -25x+9y=195 now what i would do is multiply your first equation by 3 and then your y's will cancel out and you can solve for x or you could do substitution.

OpenStudy (anonymous):

Well I am about to go to bed so I will solve for x for you then you can do the rest.. Like I said multiply the first equation by 3 all the way through 12x-9y=-117 -25+9y=195 -13x=78 x=-6 Now that you have x=6 you can solve for y by plugging 6 in for x in either equation.